I need help solving a substitution problom
-3x+y=-3
2x-5y=-11

Answer by checkley77(12844)

-3x+y=-3 or y=3x-3
2x-5y=-11
2x-5(3x-3)=-11
2x-15x+15=-11
-13x=-11-15
-13x=-26
x=-26/-13
x=2 ans.
y=3*2-3
y=6-3
y=3 ans.